# Adapting Noise to Data: Generative Flows from 1D Processes

### v2

@@ 1 Introduction
Flow-based generative models, especially score-based diffusion [32, 34], flow matching (FM) [1, 22, 23] and consistency models like the recently introduced inductive moment matching (IMM) [41], achieve state-of-the-art results in many applications. All these methods construct a probability flow from a simple latent distribution (noise) to a complex target (data) with a neural network trained to approximate this flow from limited target samples. In diffusion models, the score function directs a reverse-time SDE, while in FM, the velocity field is learned to compute trajectories via a flow ODE. Finally, consistency models like IMM learn to predict the jumps from noise to the data while factoring in the consistency of the flow trajectories. Usually, a Gaussian is used as latent distribution which causes difficulties when learning certain multimodal and heavy-tailed targets [15, 29], see Figure 5 for a heavy-tailed example. There exist only few approaches to learn the noising process, [3] fit the forward diffusion process via a learned invertible map that is trained end-to-end, [20] use metric flow matching, i.e., a neural network to adapt the path to a underlying Riemannian metric. In a related approach, [28] learns a componentwise Gaussian noise schedule, input-conditioned so each component receives its own noise level. In the setting of sampling from unnormalized target densities, [4] learn the latent noise by optimizing the mean and covariance of a Gaussian prior, while [5] learn a Gaussian mixture prior, both are trained end to end. On the other hand [25, 40] design heavy-tailed diffusions using Student-$t$ latent distributions, and [31] extend the framework to the family of $\alpha$-stable distributions.
Figure 1: FM via optimal coupling with Gaussian noise (top) and our learned noise (bottom). Samples from the latent distribution are black and from generated ones are red. The left image shows their transportation paths in green. Starting from the learned latent drastically shortened the paths. (Zoom in for the paths in the bottom image).
@@
Given any appropriate 1D process we demonstrate how to learn the componentwise neural flow by the associated conditional velocity field. We give examples besides diffusion demonstrating the flexibility of our machinery, namely the Kac process arising from the 1D damped wave equation, see [10, 16], and a process reflecting the Wasserstein gradient flow of the maximum mean discrepancy with negative distance kernel towards the uniform distribution. In contrast to diffusion, assuming a compactly supported target, these processes also have a compact support, leading to a better regularity of the corresponding velocity field. This inspired us to further adapt the process to the data and to learn the 1D noising process rather than choosing it manually. To this end, we exploit that 1D probability measures can be equivalently described by their quantile functions $Q^{i}:(0,1)\to\mathbb{R}$ which are monotone functions, and consider quantile processes $X_{t}^{i}=(1-t)X_{0}^{i}+tQ^{i}(U^{i})$, $i=1,\ldots,d$ with i.i.d. $U^{i}\sim\mathcal{U}[0,1]$ for $t\in[0,1]$. We learn the individual quantile functions $Q^{i}_{\phi}$, $i=1,\ldots,d$ such that their componentwise concatenation $\mathbf{Q}_{\phi}(\mathbf{U})\coloneqq(Q_{\phi}^{i}(U^{i}))_{i=1}^{d}$ is "close" to the data.
@@ 2.1 Absolutely Continuous Curves in Wasserstein Space
We start with a brief introduction of curves in Wasserstein spaces and basic ideas on flow matching. For more details we refer to [2] and [39]. Let $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ denote the complete metric space of probability measures with finite second moments equipped with the Wasserstein distance
@@ 2.2 Flow Matching
Unfortunately this loss function is not helpful, since we do not know the exact velocity field $v_{t}$ nor can sample from $\mu_{t}$ in the empirical expectation. However, employing the law of total probabilities, as done, e.g. in [22], we see that $\mathcal{L}(\theta)=\mathcal{L}_{\mathrm{CFM}}(\theta)+\text{const}$ with a constant not depending on $\theta$ and the Conditional Flow Matching (CFM) loss
@@ 2.3 Stochastic Processes and Velocity Fields
Consider a continuously differentiable (noising) process $(\mathbf{Y}_{t})_{t}$ with $\mathbf{Y}_{0}\equiv 0\in\mathbb{R}^{d}$ with associated velocity field $v_{t}=v^{\mathbf{Y}}_{t}(\cdot\mid 0)$ such that the pair $(\mu^{\mathbf{Y}}_{t},v^{\mathbf{Y}}_{t})$ satisfy the continuity equation (1), where $\mu^{\mathbf{Y}}_{t}$ is the law of $(\mathbf{Y}_{t})_{t}$ [^2] Existence of the velocity is given under weak assumptions by [39] Theorem 6.3.. To construct a generative model we need to create a process $(\mathbf{X}_{t})_{t}$ which can start in any sample $x_{0}$ from the target measure $\mu_{0}$. Let $\mathbf{X_{0}}\sim\mu_{0}$. Following the lines in [10], we define the mean‐reverting process by
@@
The conditional velocity field of $\mathbf{X}_{t}$ is given by (see [39, 23])
@@
Then $f(t)\coloneqq 1-t$, $g(t)\coloneqq t^{2}$ yields (independent) FM [22], and $f(t)\coloneqq\exp\left(-\frac{h(t)}{2}\right)$, $g(t)\coloneqq 1-\exp\left(-h(t)\right)$, where $h(t)\coloneqq\int_{0}^{t}\beta_{\rm min}+s(\beta_{\rm max}-\beta_{\rm min})\mathop{}\!\mathrm{d}s$ with, e.g., $\beta_{\rm min}=0.1,\,\beta_{\rm max}=20$, corresponds to processes used in score-based generative models [35], see Appendix B.
@@
Instead of considering (possibly independent) random variables $\mathbf{X}_{0},\mathbf{X}_{1}$ and their induced processes such as in (8), we can also introduce a coupling between them. Let $X_{0}\sim\mu_{0},X_{1}\sim\mu_{1}$ and consider the optimal coupling $\pi\in\Pi_{o}(\mu_{0},\mu_{1})$. Then define the induced curve $(e_{t})_{\sharp}\pi$, where $e_{t}(x,y):=(1-t)x+ty$. This curve is a geodesic between $\mu_{0}$ and $\mu_{1}$ in the Wasserstein geometry, for more details see [39]. Using the optimal coupling $\pi\in\Pi_{o}(\mu_{0},\mu_{1})$ yields an Optimal Transport FM objective
@@
The velocity field minimizing this objective then satisfies the continuity equation together with $e_{t}\sharp\pi$. In contrast to using the independent coupling, this can lead to reduced variance in training and both shorter and straighter paths, see [37, 27].
@@ 3 From One-Dimensional to Multi-Dimensional Flows
For the $d$-dimensional process $\mathbf{Y}_{t}\coloneqq(Y_{t}^{1},\dots,Y_{t}^{d})$, independence implies that its law is exactly $\mu_{t}$. Moreover, by the following proposition, the corresponding $d$-dimensional velocity field is given componentwise, see [10].
@@
In each case, the absolutely continuous curve starting in $\delta_{0}$ and the corresponding velocity field can be calculated analytically. Note that in contrast to the Wiener process $W_{t}$ usually seen in diffusion and flow matching models, the latter two processes $K_{t},U_{t}$ do not enjoy a trivial analogue in multiple dimensions: in case of $K_{t}$ the corresponding PDE (damped wave equation) is no longer mass-conserving in dimension $d\geq 3$, see [36]; in case of $U_{t}$ the mere existence of the MMD gradient flow in multiple dimensions is unclear by the lack of convexity of the MMD, see [17]. Our general construction method makes these 1D processes accessible for generative modeling in arbitrary dimensions.
@@ 3.1 Wiener Process and Diffusion Equation
However, its $L_{2}$-norm fulfills $\|{v}_{t}\|_{L_{2}(\mathbb{R},p_{t})}^{2}=\frac{d}{4t}$, and is therefore not integrable over time, i.e. $\|{v}_{t}\|_{L_{2}(\mathbb{R},p_{t})}\notin L_{2}(0,1)$. In practice, instability issues caused by this explosion at times close to the target need to be avoided by e.g. time truncations, see e.g. [21]. For a heuristic analysis also including drift-diffusion flows, we refer to [26]. Note that in the case of diffusion, there is no significant distinction between the uni- and multivariate setting. Figure 3.2 shows the generation of samples from a weighted Gaussian Mixture Model (GMM) using Flow Matching and the Wiener process as our noising process. As described in Section 2.3 we define the mean reverting process and use schedules f(t)=1tf(t)=1-t and $g(t)=t^{2}$.
@@ 3.2 Kac Process and Damped Wave Equation
The Kac process was recently used in generative modeling by some of the authors in [10]. It originates from a discrete random walk, which starts in 0 and moves with velocity parameter $c>0$ in one direction until it reverses its direction with probability $a\Delta_{t}$, $a>0$, see [19]. A continuous-time analogue is given by the Kac process which is defined using the homogeneous Poisson point process $N_{t}$ with rate $a$, i.e. i) $N_{0}=0$; ii) the increments of $N_{t}$ are independent, iii) $N_{t}-N_{s}\sim{\rm Poi}\big(a(t-s)\big)$ for all $0\leq s<t$. Now the Kac process starting in 0 is given by
@@
By the following proposition, the Kac process is related to the damped wave equation, also known as telegrapher’s equation, and its probability distribution admits a computable vector field such that the continuity equation is fulfilled. For a proof we refer to our paper [10].
@@
Interestingly, the damped wave equation (20) is closely related to the diffusion equation via Kac’ insertion method. In particular, diffusion can be seen as "an infinitely $a$-damped wave with infinite propagation speed cc". Note that the diffusion-related concept of particles traveling with infinite speed violates Einstein’s laws of relativity and has therefore found resistance in the physics community [6, 8, 38, 36]. Figure 3.2 shows the generation of samples from a weighted Gaussian Mixture Model (GMM) using Flow Matching and the Kac process as our noising process. As described in Section 2.3 we define the mean reverting process and use schedules f(t)=1tf(t)=1-t and $g(t)=t^{2}$.
A generated trajectory from a Flow Matching model trained using the conditional density and velocity given by the Kac process with $(a,c)=(9,3)$.
@@ 3.3 Uniform Process and Gradient Flow of MMD Functional
stands out for its flexible flow behavior between distributions of different support [17]. We have have the following proposition which proof can be found in Appendix A. Note that the proof is already based on so-called quantile function, which we consider in the next section.
@@
Figure 3.2 shows the generation of samples from a weighted Gaussian Mixture Model (GMM) using Flow Matching and the process induced by the MMD gradient flow as our noising process. As described in Section 2.3 we define the mean reverting process and use schedules f(t)=1tf(t)=1-t and $g(t)=t$.
@@ 3.4 Scaled Latent Distributions
Inspired by [25] in Figure 5 we highlight how different choices of latent $Z$ can heavily affect the sampling performance. For latent distributions without (heavy) tails, the FM model fails to capture the tails of the data distribution. This motivates us to learn a data-adapted process $Y_{t}:=tZ$ by learning the terminal distribution $Z$.
Figure 6: Sampling of Neal’s funnel with different latent distributions [^5] Note that we used the independent coupling for training of these models and we pretrained our learned latent (instead of training jointly). We also used z-score normalization.. From left to right with uniform ([1,1][-1,1]), standard Gaussian, Student-T (with parameters $(20,4)$ inspired by the choice in [25]) and our learned distribution. The last two heavy-tailed noises perform significantly better.
@@ 4.1 Quantile Processes and Interpolants
The proof is given Appendix C. Proposition 6 allows us to also apply the concept of consistency models to our quantile process (36). The shared idea of these models is to predict the jumps from the process $Z_{t}$ to the target $X_{0}$, while factoring in the consistency of the trajectory of $Z_{t}$ via $Z_{s}$, $0<s<t$. In FM, this consistency of the flow is usually neglected as only single points on the FM paths are sampled. Also, consistency models as one-step or multistep samplers usually are in no need of velocity fields. In the Appendix C, we demonstrate by means of the recently proposed inductive moment matching (IMM) [41], that our formulation via quantile interpolants fits seamlessly into the consistency framework.

### v3

@@ 1 Introduction
Flow-based generative models, especially score-based diffusion [pmlr-v37-sohl-dickstein15, SE2019], flow matching (FM) [albergo2023stochastic, lipman2023flow, liu2022rectified] and consistency models like the recently introduced inductive moment matching (IMM) [IMM2025], achieve state-of-the-art results in many applications. All these methods construct a probability flow from a simple latent distribution (noise) to a complex target (data) with a neural network trained to approximate this flow from limited target samples. In diffusion models, the score function directs a reverse-time SDE, while in FM, the velocity field is learned to compute trajectories via a flow ODE. Finally, consistency models like IMM learn to predict the jumps from noise to the data while factoring in the consistency of the flow trajectories. Usually, a Gaussian is used as latent distribution which causes difficulties when learning certain multimodal and heavy-tailed targets [HN2021, SBDD2022], see Figure 5 for a heavy-tailed example. There exist only few approaches to learn the noising process, [bartosh2025neuralflowdiffusionmodels] fit the forward diffusion process via a learned invertible map that is trained end-to-end, [kapusniak2024metric] use metric flow matching, i.e., a neural network to adapt the path to a underlying Riemannian metric. In a related approach, [sahoo2024mulan] learns a componentwise Gaussian noise schedule, input-conditioned so each component receives its own noise level. In the setting of sampling from unnormalized target densities, [Blessing2025UDB] learn the latent noise by optimizing the mean and covariance of a Gaussian prior, while [blessing2025endtoendlearninggaussianmixture] learn a Gaussian mixture prior, both are trained end to end. On the other hand [pandey2024heavytaileddiffusionmodels, zhang2024longtailed] design heavy-tailed diffusions using Student-$t$ latent distributions, and [HTDL2025] extend the framework to the family of $\alpha$-stable distributions.
Figure 1: FM via optimal coupling with Gaussian noise (left) and our learned noise (right). Samples from the latent distribution are black and from generated ones are red. The left image shows their transportation paths in green. Starting from the learned latent drastically shortened the paths. (Zoom in for the paths in the bottom image).
@@
Given any appropriate 1D process we demonstrate how to learn the componentwise neural flow by the associated conditional velocity field. We give examples besides diffusion demonstrating the flexibility of our machinery, namely the Kac process arising from the 1D damped wave equation, see [DCFS2025, distillkac2025], and a process reflecting the Wasserstein gradient flow of the maximum mean discrepancy with negative distance kernel towards the uniform distribution. In contrast to diffusion, assuming a compactly supported target, these processes also have a compact support, leading to a better regularity of the corresponding velocity field. This inspired us to further adapt the process to the data and to learn the 1D noising process rather than choosing it manually. To this end, we exploit that 1D probability measures can be equivalently described by their quantile functions $Q^{i}:(0,1)\to\mathbb{R}$ which are monotone functions, and consider quantile processes $X_{t}^{i}=(1-t)X_{0}^{i}+tQ^{i}(U^{i})$, $i=1,\ldots,d$ with i.i.d. $U^{i}\sim\mathcal{U}[0,1]$ for $t\in[0,1]$. We learn the individual quantile functions $Q^{i}_{\phi}$, $i=1,\ldots,d$ such that their componentwise concatenation $\mathbf{Q}_{\phi}(\mathbf{U})\coloneqq(Q_{\phi}^{i}(U^{i}))_{i=1}^{d}$ is "close" to the data.
@@ 2.1 Absolutely Continuous Curves in Wasserstein Space
We start with a brief introduction of curves in Wasserstein spaces and basic ideas on flow matching. For more details we refer to [BookAmGiSa05] and [WS2025]. Let $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ denote the complete metric space of probability measures with finite second moments equipped with the Wasserstein distance
@@ 2.2 Flow Matching
Unfortunately this loss function is not helpful, since we do not know the exact velocity field $v_{t}$ nor can sample from $\mu_{t}$ in the empirical expectation. However, employing the law of total probabilities, as done, e.g. in [lipman2023flow], we see that $\mathcal{L}(\theta)=\mathcal{L}_{\mathrm{CFM}}(\theta)+\text{const}$ with a constant not depending on $\theta$ and the Conditional Flow Matching (CFM) loss
@@ 2.3 Stochastic Processes and Velocity Fields
Consider a continuously differentiable (noising) process $(\mathbf{Y}_{t})_{t}$ with $\mathbf{Y}_{0}\equiv 0\in\mathbb{R}^{d}$ with associated velocity field $v_{t}=v^{\mathbf{Y}}_{t}(\cdot\mid 0)$ such that the pair $(\mu^{\mathbf{Y}}_{t},v^{\mathbf{Y}}_{t})$ satisfy the continuity equation (1), where $\mu^{\mathbf{Y}}_{t}$ is the law of $(\mathbf{Y}_{t})_{t}$ [^2] Existence of the velocity is given under weak assumptions by [WS2025] Theorem 6.3.. To construct a generative model we need to create a process $(\mathbf{X}_{t})_{t}$ which can start in any sample $x_{0}$ from the target measure $\mu_{0}$. Let $\mathbf{X_{0}}\sim\mu_{0}$. Following the lines in [DCFS2025], we define the mean‐reverting process by
@@
The conditional velocity field of $\mathbf{X}_{t}$ is given by (see [WS2025, liu2022rectified])
@@
Then $f(t)\coloneqq 1-t$, $g(t)\coloneqq t^{2}$ yields (independent) FM [lipman2023flow], and $f(t)\coloneqq\exp\left(-\frac{h(t)}{2}\right)$, $g(t)\coloneqq 1-\exp\left(-h(t)\right)$, where $h(t)\coloneqq\int_{0}^{t}\beta_{\rm min}+s(\beta_{\rm max}-\beta_{\rm min})\mathop{}\!\mathrm{d}s$ with, e.g., $\beta_{\rm min}=0.1,\,\beta_{\rm max}=20$, corresponds to processes used in score-based generative models [song2021scorebasedgenerativemodelingstochastic], see Appendix B.
@@
Instead of considering (possibly independent) random variables $\mathbf{X}_{0},\mathbf{X}_{1}$ and their induced processes such as in (8), we can also introduce a coupling between them. Let $X_{0}\sim\mu_{0},X_{1}\sim\mu_{1}$ and consider the optimal coupling $\pi\in\Pi_{o}(\mu_{0},\mu_{1})$. Then define the induced curve $(e_{t})_{\sharp}\pi$, where $e_{t}(x,y):=(1-t)x+ty$. This curve is a geodesic between $\mu_{0}$ and $\mu_{1}$ in the Wasserstein geometry, for more details see [WS2025]. Using the optimal coupling $\pi\in\Pi_{o}(\mu_{0},\mu_{1})$ yields an Optimal Transport FM objective
@@
The velocity field minimizing this objective then satisfies the continuity equation together with $e_{t}\sharp\pi$. In contrast to using the independent coupling, this can lead to reduced variance in training and both shorter and straighter paths, see [tong2024improvinggeneralizingflowbasedgenerative, pooladian2023multisampleflowmatchingstraightening].
@@ 3 From One-Dimensional to Multi-Dimensional Flows
For the $d$-dimensional process $\mathbf{Y}_{t}\coloneqq(Y_{t}^{1},\dots,Y_{t}^{d})$, independence implies that its law is exactly $\mu_{t}$. Moreover, by the following proposition, the corresponding $d$-dimensional velocity field is given componentwise, see [DCFS2025].
@@
In each case, the absolutely continuous curve starting in $\delta_{0}$ and the corresponding velocity field can be calculated analytically. Note that in contrast to the Wiener process $W_{t}$ usually seen in diffusion and flow matching models, the latter two processes $K_{t},U_{t}$ do not enjoy a trivial analogue in multiple dimensions: in case of $K_{t}$ the corresponding PDE (damped wave equation) is no longer mass-conserving in dimension $d\geq 3$, see [TL2016]; in case of $U_{t}$ the mere existence of the MMD gradient flow in multiple dimensions is unclear by the lack of convexity of the MMD, see [HGBS2022]. Our general construction method makes these 1D processes accessible for generative modeling in arbitrary dimensions.
@@ 3.1 Wiener Process and Diffusion Equation
However, its $L_{2}$-norm fulfills $\|{v}_{t}\|_{L_{2}(\mathbb{R},p_{t})}^{2}=\frac{d}{4t}$, and is therefore not integrable over time, i.e. $\|{v}_{t}\|_{L_{2}(\mathbb{R},p_{t})}\notin L_{2}(0,1)$. In practice, instability issues caused by this explosion at times close to the target need to be avoided by e.g. time truncations, see e.g. [soft2021]. For a heuristic analysis also including drift-diffusion flows, we refer to [P2022]. Note that in the case of diffusion, there is no significant distinction between the uni- and multivariate setting. Figure 5 shows the generation of samples from a weighted Gaussian Mixture Model (GMM) using Flow Matching and the Wiener process as our noising process. As described in Section 2.3 we define the mean reverting process and use schedules f(t)=1tf(t)=1-t and $g(t)=t^{2}$.
@@ 3.2 Kac Process and Damped Wave Equation
The Kac process was recently used in generative modeling by some of the authors in [DCFS2025]. It originates from a discrete random walk, which starts in 0 and moves with velocity parameter $c>0$ in one direction until it reverses its direction with probability $a\Delta_{t}$, $a>0$, see [KAC1974]. A continuous-time analogue is given by the Kac process which is defined using the homogeneous Poisson point process $N_{t}$ with rate $a$, i.e. i) $N_{0}=0$; ii) the increments of $N_{t}$ are independent, iii) $N_{t}-N_{s}\sim{\rm Poi}\big(a(t-s)\big)$ for all $0\leq s<t$. Now the Kac process starting in 0 is given by
@@
By the following proposition, the Kac process is related to the damped wave equation, also known as telegrapher’s equation, and its probability distribution admits a computable vector field such that the continuity equation is fulfilled. For a proof we refer to our paper [DCFS2025].
@@
Interestingly, the damped wave equation (20) is closely related to the diffusion equation via Kac’ insertion method. In particular, diffusion can be seen as "an infinitely $a$-damped wave with infinite propagation speed cc". Note that the diffusion-related concept of particles traveling with infinite speed violates Einstein’s laws of relativity and has therefore found resistance in the physics community [C1958, C1963, V1958, TL2016]. Figure 5 shows the generation of samples from a weighted Gaussian Mixture Model (GMM) using Flow Matching and the Kac process as our noising process. As described in Section 2.3 we define the mean reverting process and use schedules f(t)=1tf(t)=1-t and $g(t)=t^{2}$.
Figure 5: A generated trajectory from a Flow Matching model trained using the conditional density and velocity given by the Kac process with $(a,c)=(9,3)$.
@@ 3.3 Uniform Process and Gradient Flow of MMD Functional
stands out for its flexible flow behavior between distributions of different support [HGBS2022]. We have have the following proposition which proof can be found in Appendix A. Note that the proof is already based on so-called quantile function, which we consider in the next section.
@@
Figure 5 shows the generation of samples from a weighted Gaussian Mixture Model (GMM) using Flow Matching and the process induced by the MMD gradient flow as our noising process. As described in Section 2.3 we define the mean reverting process and use schedules f(t)=1tf(t)=1-t and $g(t)=t$.
@@ 3.4 Scaled Latent Distributions
Inspired by [pandey2024heavytaileddiffusionmodels] in Figure 5 we highlight how different choices of latent $Z$ can heavily affect the sampling performance. For latent distributions without (heavy) tails, the FM model fails to capture the tails of the data distribution. This motivates us to learn a data-adapted process $Y_{t}:=tZ$ by learning the terminal distribution $Z$.
Figure 7: Sampling of Neal’s funnel with different latent distributions [^5] Note that we used the independent coupling for training of these models and we pretrained our learned latent (instead of training jointly). We also used z-score normalization.. From left to right with uniform ([1,1][-1,1]), standard Gaussian, Student-T (with parameters $(20,4)$ inspired by the choice in [pandey2024heavytaileddiffusionmodels]) and our learned distribution. The last two heavy-tailed noises perform significantly better.
@@ 4.1 Quantile Processes and Interpolants
The proof is given Appendix C. Proposition 6 allows us to also apply the concept of consistency models to our quantile process (36). The shared idea of these models is to predict the jumps from the process $Z_{t}$ to the target $X_{0}$, while factoring in the consistency of the trajectory of $Z_{t}$ via $Z_{s}$, $0<s<t$. In FM, this consistency of the flow is usually neglected as only single points on the FM paths are sampled. Also, consistency models as one-step or multistep samplers usually are in no need of velocity fields. In the Appendix C, we demonstrate by means of the recently proposed inductive moment matching (IMM) [IMM2025], that our formulation via quantile interpolants fits seamlessly into the consistency framework.